OpenStudy (anonymous):

[8.07] Factor Completely: 25x2 − 49 (25x + 7)(x − 7) (25x - 7)(x − 7) (5x + 7)(5x − 7) (5x − 7)(5x − 7)

jimthompson5910 (jim_thompson5910):

hint: x^2 - y^2 = (x - y)(x + y)

OpenStudy (anonymous):

Ok lemme work this out, or at least try

jimthompson5910 (jim_thompson5910):

ok tell me what you get

jimthompson5910 (jim_thompson5910):

it may help to recognize that 25x^2 = (5x)^2 49 = 7^2